OpenStudy (anonymous):

Using pie = 22/7. How do I find the area of the circle of 42ft?

Parth (parthkohli):

I like apple pie.

Parth (parthkohli):

I wonder if these schools are reliable; they spell arguably the most important constant wrong.

OpenStudy (anonymous):

god i hate math teachers. now students will think \(\pi=\frac{22}{7}\) a rather silly approximation based on basically nothing

OpenStudy (shubhamsrg):

pie could also be interpreted as (pi)*(e)

OpenStudy (anonymous):

lol

Parth (parthkohli):

Or p * i * e.

OpenStudy (anonymous):

\(A=\pi r^2\) so \(A=\pi \times 42^2\)

OpenStudy (shubhamsrg):

Unit being (feet)^2

Parth (parthkohli):

42 can be the diameter too.

OpenStudy (theviper):

According to question 42ft is circumference ?

OpenStudy (shubhamsrg):

Yeah, should be assumed circumference according to the language.

OpenStudy (anonymous):

Forgive me for the wrong abbreviations. In the book 42ft was changed to (21)2, why?

OpenStudy (theviper):

\[2\pi r=42\ ft\]\[=>r=42/2\pi\] Therefore area=\(\LARGE{\pi r^2=\pi(\frac{42}{2\pi})^2}=?\)

OpenStudy (theviper):

???????????